4. Find the LaPlace transforms for the following functions. (5pts) (a) f(t) = 3t2 - sin 3t (b) t +1 (2-et ostsil t21
Answer #1

(a) f(1) = 312-sinat Laplace transfoom of t, DE N (natural Durobers) h! Itos Laplace transforms of Siscat) - b sta² LCfCt))ostal (b) f(t) = t+1 {2-et ta! f(t) = +1 +(2-et-(t+1)) u (t-1) ++1 + [li-et- t)u(t-1)] LCFC)] = L(+)+L(1)+461-et)u(t-1)] F(tFu) - L[f1t)) = -14[] -- [] + 1 + S s? - ersto But L[fct>] =l[rct-a)ult-a)] = 8°S F(t) -S + ㅗ sa = [0–1)n(7-72-1937 :: e(5+1)